“I know headquarters wants us to add that new product line,” said Dell Havasi, manager of Billings Company’s Office Products Division. “But I want to see the numbers before I make any move. Our division’s return on investment (ROI) has led the company for three years, and I don’t want any letdown.”

Billings Company is a decentralized wholesaler with five autonomous divisions. The divisions are evaluated on the basis of ROI, with year-end bonuses given to the divisional managers who have the highest ROIs. Operating results for the company’s Office Products Division for this year are given below:

Sales

$

22,700,000

Variable expenses

14,363,700

Contribution margin

8,336,300

Fixed expenses

6,175,000

Net operating income

$

2,161,300

Divisional average operating assets

$

5,675,000

The company had an overall return on investment (ROI) of 16.00% this year (considering all divisions). Next year the Office Products Division has an opportunity to add a new product line that would require an additional investment that would increase average operating assets by $3,938,000. The cost and revenue characteristics of the new product line per year would be:

Sales

$9,800,000

Variable expenses

65% of sales

Fixed expenses

$2,582,900

Required:

1. Compute the Office Products Division’s ROI for this year.

2. Compute the Office Products Division’s ROI for the new product line by itself.

3. Compute the Office Products Division’s ROI for next year assuming that it performs the same as this year and adds the new product line.

4. If you were in Dell Havasi’s position, would you accept or reject the new product line?

5. Why do you suppose headquarters is anxious for the Office Products Division to add the new product line?

6. Suppose that the company’s minimum required rate of return on operating assets is 13% and that performance is evaluated using residual income.

a. Compute the Office Products Division’s residual income for this year.

b. Compute the Office Products Division’s residual income for the new product line by itself.

c. Compute the Office Products Division’s residual income for next year assuming that it performs the same as this year and adds the new product line.

d. Using the residual income approach, if you were in Dell Havasi’s position, would you accept or reject the new product line?

Req1-3

1.

ROI for this year

38.08

%

2.

ROI for the new product line by itself

21.51

%

3.

ROI for next year

31.30

%

Req4

If you were in Dell Havasi’s position, would you accept or reject the new product line?

Acceptradio button checked1 of 2

Rejectradio button unchecked2 of 2

Req5

Why do you suppose headquarters is anxious for the Office Products Division to add the new product line?

Adding the new line would increase the company's overall ROI.radio button checked1 of 2

Adding the new line would decrease the company's overall ROI.radio button unchecked2 of 2

Req6A-6C


1.

Residual income for this year

2.

Residual income for the new product line by itself

3.

Residual income for next year

Using the residual income approach, if you were in Dell Havasi’s position, would you accept or reject the new product line?

Acceptradio button unchecked1 of 2

Rejectradio button unchecked2 of 2
Answer #1

1.ROI = Net Operating Income/Average Assets

= 2,161,300/5,675,000

= 38.08%

2.ROI of new product line = (9,800,000*35%-2,582,900)/3,938,000

= 21.51%

3.ROI when new line is added

= (2,161,300+ 847,100)/9,613,000

= 31.29%

4.Reject, since will reduce ROI

5.Because ROI of new line is higher than the overall ROI of the company

4.Residual Income = Net operating Income – Average operating assets*Minimum required rate of return

a.Office Products residual income = 2,161,300 – 5,675,000*13%

= $1,423,550

b.New Product Line added = 847,100– 3,938,000*13%

= $335,160

c.Residual Income = 1,423,550+335,160 = $1,758,710

d.Accept as Residual Income has increased